OpenStudy (anonymous):

Algebra 2 help fan and medal help me plzzz

OpenStudy (anonymous):

2) Using both fundamental Theorem and Descartes’ rule of signs, prove to the construction foreman that your function matches your graph. Use complete sentences.

OpenStudy (anonymous):

My equation x^3+3x^2-26x-84

OpenStudy (anonymous):

@perl @Samus_Aran

OpenStudy (perl):

i would need to see the entire question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The rollercoaster will run from right to left, so work in that order when selecting points. 1st point: 6 2nd point: -2 3rd point: -7 1) Using the points above as zeros, construct the polynomial function, f(x), that will be the pasth of your rollercoaster. Show all of your work. 2) Using both fundamental Theorem and Descartes’ rule of signs, prove to the construction foreman that your function matches your graph. Use complete sentences. 3) Solve for the y-intercept for your function, f(x), and then construct a rough graph of your rollercoaster. If your y-intercept is off the graph, give the coordinates of the y-intercept.

OpenStudy (anonymous):

i already did 1 and 3 i just need help with two and the ? are just there they just appeared when i copied and pasted them

OpenStudy (perl):

hmm

OpenStudy (perl):

can you upload the original document or worksheet

OpenStudy (anonymous):

ummm yeah

OpenStudy (anonymous):

thats it with what my answers are so far

OpenStudy (perl):

ok one sec

OpenStudy (perl):

here is an older response http://openstudy.com/study#/updates/532476f8e4b087b3d40f8a72

OpenStudy (anonymous):

Here is one I found https://answers.yahoo.com/question/index?qid=20140620131439AAZMNJQ

OpenStudy (anonymous):

i looked at those and they didnt apply to my equatiuon so i was confused

OpenStudy (perl):

given the zeros, you can give the general form of polynomial f(x)=a(x+7)(x+2)(x-6)

OpenStudy (anonymous):

ok lol idk what the 2 rules are so could you explain the question to me plz

OpenStudy (perl):

ok

OpenStudy (perl):

we want to prove that your function matches the graph of the rollercoaster

OpenStudy (anonymous):

yeah

OpenStudy (perl):

you have a problem with your function. I got y=(x+7)(x+2)(x-6) = x^3+3 x^2-40 x-84

OpenStudy (perl):

(x+7)(x+2)(x-6) = =(x^2 + 7x + 2x + 14) ( x - 6) =( x^2 + 9x + 14) ( x - 6) = x^3 + 9x^2 + 14x - 6x^2 -54x - 84 = x^3 + 3x^2 - 40x - 84

OpenStudy (anonymous):

hmmm

OpenStudy (perl):

you made a mistake somewhere

OpenStudy (anonymous):

i guess lol

OpenStudy (anonymous):

so what next

OpenStudy (perl):

also I would multiply by a factor to vertically compress or shrink it

OpenStudy (anonymous):

ok so wb for question 2

OpenStudy (perl):

so what does wb stand for

OpenStudy (anonymous):

what about

OpenStudy (perl):

ok

OpenStudy (perl):

so you want to multiply your function by 1/20

OpenStudy (perl):

y = 1/20 ( x^3 + 3x^2 - 40x - 84)

OpenStudy (anonymous):

for number 2??

OpenStudy (perl):

still for number 1

OpenStudy (perl):

im looking for a number that goes nicely into 84 and those numbers

OpenStudy (anonymous):

i already did number one lol can you help me with number 2 plzzzz

OpenStudy (anonymous):

@shereenkhan

OpenStudy (perl):

your graph is a bit extreme

OpenStudy (anonymous):

ik i changed it

OpenStudy (perl):

ok

OpenStudy (perl):

i mean, it would be too steep

OpenStudy (anonymous):

yeah lol i fixed it

OpenStudy (anonymous):

@gorv

OpenStudy (perl):

First of all, the Fundamental Theorem of Algebra guarantees that f(x) has at least one real root (one real x intercept )

OpenStudy (anonymous):

ok

OpenStudy (perl):

we're not really going to use the FTA much here (fund. theorem of algebra), but you can mention that the FTA guarantees 3 complex roots (imaginary roots come in pairs, so you have to have one real root )

OpenStudy (perl):

the descartes rule of signs , we look at f(x) and f(-x), and count the number of sign changes

OpenStudy (perl):

f(x) = x^3 + 3x^2 - 40x - 84 , this has one sign change

OpenStudy (gorv):

f(x)= x^3+3x^2-26x-84 we check the sign change happen from left to right here is only one sign change + to -

OpenStudy (gorv):

now for f(-x)=-x^3+3x^2+26x-84 here 2 sign change are there

OpenStudy (anonymous):

ok idk what to write for the answer ecause im noy sure what its asking

OpenStudy (anonymous):

can you give me an example of hat youd right plz???

OpenStudy (perl):

it goes from + to - , just once. so the number of positive real zeros are 1 or less by a multiple of 2. but you cant subtrat 1 by a multiple of 2 , you will get -1 , -3 , -5 , etc

OpenStudy (anonymous):

what lol im so confused

OpenStudy (gorv):

we are given point 6 , -7 and -2 f(x) = (x - 6)(x - (-2))(x - (-7)) = (x - 6)(x + 2)(x + 7) = (x - 6)(x^2 + 9x + 14) = x(x^2 + 9x + 14) - 6(x^2 + 9x + 14) = x^3 + 3x^2 - 40x - 84. so as it is of degree 3 so it will have 3 zeros As for Descartes' Rule of signs: Since f(x) has 1 sign change, f has exactly 1 positive zero. Moreover, f(-x) = -x^3 + 3x^2 + 40x - 84 also has 2 sign changes. So, f has either 2 or 0 negative zeros. as f(-2) = 0 and f(-7) = 0, we see that f indeed has exactly 2 negative zeros.)

OpenStudy (anonymous):

ok ty!! so wb fta

OpenStudy (perl):

the fta just says there are possible at most 3 real zeroes ,

OpenStudy (perl):

3 x intercepts

OpenStudy (anonymous):

ok so how do i write that in an answer??

OpenStudy (perl):

this is a good explanation of descartes rule of signs http://mathworld.wolfram.com/DescartesSignRule.html

OpenStudy (perl):

you can mention that since the polynomial has degree 3 (highest exponent is 3) we will have a maximum of 3 x intercepts (we can't have more than 3 x intercepts, but we could have less then 3 )

OpenStudy (perl):

the FTA gives us some information about the graph, just by the degree . but to narrow it down further we use Descartes law of signs

OpenStudy (perl):

so in other words, the foreman is skeptical that your polynomial function accurately graphs the rollercoaster. the foreman might wonder, what if your function has more than three x-intercepts , maybe it has 4 or 5. you tell him, thats not possible by the FTA

OpenStudy (perl):

you dont really have to say much about the FTA, just one sentence.